OpenStudy (anonymous):

APY=(1+p)^(n)-1 <--- Is that the right formula Bridgette opened a store credit card to purchase a business suit for $463. She put the entire purchase on the credit card. Her APR is 14.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months? $396.97 $412.83 $397.35 $445.63

OpenStudy (tkhunny):

Not really, on the formula. It might be close, but it's not clear where you are going. This one is trickier than most, because the payments are not level. Just work it out one payment at a time and see where it leads. i = 0.1499 j = 0.1499/12 = 0.0124916667 At Purchase: 463.00 Al month 1: 463.00(1+j) - 0.05*463.00(1+j) You need to understand where all those pieces came from. Give it some thought and work out the new balance at Month 1.

OpenStudy (anonymous):

I got 445.344207 ? :D

OpenStudy (tkhunny):

You should use more decimal places in your intermediate calculations. 445.344459 However, it comes out 445.34, anyway. Now, let's think about it a little. Can we simplify our lives? i = 0.1499 j = 0.1499/12 = 0.0124916667 At Purchase: 463.00 At Month 1: 463.00(1+j) - 0.05*463.00(1+j) A little algebra shows another useful expression for the value at Month 1 At Month 1: 463.00(1+j)(1 - 0.05) = 463.00*0.95*(1+j) Convince yourself that these expressions provide the same result.